Electronic Distribution has a defined benefit pension plan. Characteristics of the plan during 2018 are as follows: ($ millions) PBO balance, January 1 $ 520 Plan assets balance, January 1 250 Service cost 95 Interest cost 65 Gain from change in actuarial assumption 34 Benefits paid (44) Actual return on plan assets 17 Contributions 2018 85 The expected long-term rate of return on plan assets was 8%. There were no AOCI balances related to pensions on January 1, 2018, but at the end of 2018, the company amended the pension formula creating a prior service cost of $17 million. (Enter your answers in millions (i.e., 10,000,000 should be entered as 10).) Required: 1. Calculate the pension expense for 2018. 2. Prepare the journal entry to record pension expense, gains or losses, prior service cost, funding, and payment of benefits for 2018. 3. What amount will Electronic Distribution report in its 2018 balance sheet as a net pension asset or net pension liability?
Answer #1

1.

Calculate the pension expense as follows:

Pension expense = Service cost + interest cost - expected return on the plan assets + amortization of prior service cost

Pension expense = 95 + 65 - 20 +0

Pension expense = $ 140 million

2.

Prepare the journal entries as follows:

No Accounts Debit Credit
1 Pension expense 140
Plan assets (250 * 8%) 20
Projected benefit obligation 160
(Recording of pension expense)
2 Other comprehensive income 17
Projected benefit obligation 17
(Recording of prior service cost)
3 Projected benefit obligation 34
Other comprehensive income - Gain 34
(Recording gain)
4 Other comprehensive income - Loss (17-20) 3
Plan assets 3
(Recording loss)
5 Plan assets 85
Cash 85
(Record contributions )
6 Projected benefit obligation 44
Plan assets 44
(Record benefits paid)

3.

Determine the amount of PBO, closing as follows:

PBO, Closing = Opening balance + service cost + interest cost - gain from change in actuarial assumption + prior service cost - benefits paid

PBO, Closing = 520 + 95 + 65 - 34 + 17 - 44

PBO, Closing = $ 619 million

Determine the amount of closing plant asset as follows:

Closing plant asset = opening balance + Actual return on plan assets + Contributions 2018 - benefits paid

Closing plant asset = 250 + 17 + 85 - 44

Closing plant asset = $ 308 million

Determine the net pension liability as follows:

Net pension liability = 619 - 308

Net pension liability = $ 311 million